# 25mm25mm30mm30mm

### v1

@@ Abstract.
The construction of Sweedler duality is an important tool in the theory of Hopf algebras over a field, which is a right adjoint to the dual algebra functor. In this paper, we study the Sweedler duality of Hom-algebras and their Hom-modules. We delve into the structure of Hom-coalgebras and derive the linear morphisms associated with them. Additionally, as an application, we present the (right) Hom-(co)module morphisms under the Sweedler duality.
@@
Primary 17A30; Secondary 17A60, 17D30
@@ 1. Introduction
Algebraic deformation theory was first studied systematically and in depth by Gerstenhaber [5] and is now one of the most important branches of algebra. In recent years, the introduction of another type of algebraic deformation algebra, namely Hom-algebra, has attracted the attention of many algebraists. In 2006, Hartwing, Larsson and Silvestrov [7] introduced Hom-Lie algebras to describe Q-transformations in Witt and Virasoro algebras. In 2008, Makhlouf and Silvestrov [10] extended the structure of Hom-Lie algebras to quasi-Hom-Lie algebras and quasi-Lie algebras where they provided a different method for constructing Hom-Lie algebras, which is based on the process of generalising the construction of Lie algebras from associative algebras by defining the Lie bracket as a commutator. Building on this foundation, they [11][15] generalized the general (co)algebraic structure to Hom-(co)algebras which essentially extends the concept of associative algebras by deforming the associativity law into an $\alpha$-associativity condition. Later, Yau [17][18][19][21] proposed the concepts of Hom-(co)modules, Hom-Hopf modules, and the duality of Hom-algebras, thus enriching the field of Hom-algebra theory. As an emerging field within algebra, (Bi)Hom-algebra is intimately connected to category theory (See,e.g.,[22][14]), combinatorics (See,e.g.,[12]), braided group representations (See,e.g.,[9]), and quantum algebras (See,e.g.,[13]), producing a wealth of research outcomes in recent years.
@@
The exponentials of infinitesimal characters and their convolutions are known to produce characters with elegant combinatorial counterparts [2]. In a Hopf algebra, the set of characters forms a group that acquires a natural structure of an (infinite-dimensional) Lie group under specific conditions, which are commonly met in practice, such as grading. All these elements pertain to the largest dual one can consider for a Hopf algebra, known as Sweedler duality. This duality is, in a sense, connected to finite state automaton theory [3] and algebraic combinatorial theory [16].
@@
In the finite dimensional picture, the structure of the Hom-coalgebra is obtained from the Hom-algebra by an universal dual approach, and such a conclusion is known. This naturally leads to the question that whether the above structure can also be achieved in the infinite dimensional case? In other words, is it possible to realize the induction of Hom-algebras to Hom-coalgebras (resp. Hom-modules to Hom-comodules) by constructing a duality under infinite dimensions (as we called it above the Sweedler duality)?
@@
In this paper, we investigate the Sweedler duality of Hom-algebras and their induction of Hom-coalgebras in general. We present the new structure of the Hom-coalgebra that is induced by the Hom-algebra within the context of Sweedler duality. Moreover, we extend our findings to Hom-modules and establish the connection between Hom-modules and Hom-comodules.
@@
The layout of this paper is as follows. In Section 2, we review some basic concepts and lemmas that discuss the relationship between Hom-algebra and Hom-coalgebra. The structure of the Hom-coalgebra under Sweedler duality is obtained in Theorem 2. A morphism of Hom-coalgebra is derived using the condition above in Corollary 2.10. In Section 3, we present a right Hom-comodule of Hom-coalgebra, which is induced by the right-module of Hom-algebra under Sweedler duality, as shown in Theorem 4. The (right) Hom-comodule morphism is also provided in Theorem 5.
Throughout this paper, for the mapping role of elements, we will note them like this $f(v):=\langle f,v\rangle$ for all $f\in G^{*}$, $v\in G$. More details about notations can be found in [5]. Throughout, we will assume that the characteristic of the field $\mathbb{K}$ is equal to zero.
2. The Sweedler duality for Hom-algebras
@@
In this section, we will first provide a brief overview of the fundamental concepts related to the structures of Hom-(co)algebra and the interrelationship between Hom-(co)algebras (see Definition 2.3 in [20]). Subsequently, we introduce the concept of Sweedler duality, which differs from the conventional duality (see Chapter 6 in [5]).
@@
A Hom-algebra is a triple $(G,\mu,\alpha)$, where $G$ is a $\mathbb{K}$-linear space, $\mu:G\otimes G\rightarrow G$ is a bilinear multiplication with $\mu(g\otimes g^{\prime})=gg^{\prime}$. For simplicity when there is no confusion, we will omit the sign $\mu$. $\alpha:G\rightarrow G$ is a $\mathbb{K}$-linear space homomorphism. Otherwise it satisfies the following conditions
@@
In this context, we will refer to a Hom-algebra $(G,\mu,\alpha)$ as $G$. Let us consider two Hom-algebras, $(G,\mu,\alpha)$ and $(G^{\prime},\mu^{\prime},\alpha^{\prime})$, both defined over a field $\mathbb{K}$. Let $f:G\to G^{\prime}$ be a $\mathbb{K}$-linear map. The linear map $f$ is said to be a morphism of Hom-algebras if it satisfies the following conditions:
@@
In other words, $f$ is a morphism of Hom-algebras if it commutes with the multiplication operations and respects the Hom-structures of $G$ and $G^{\prime}$.
@@
Let $(H,\mu,\alpha)$ be a subspace of $G$. A Hom-subalgebra of G is defined as a subspace of G that satisfies the following conditions:
$bc\in H,\quad\alpha(b)\in H,$
for any $b,c\in H$. In particular, if $ab\in H$ and $ba\in H$, for any $a\in G$ and $b\in H$. The Hom-algebra $G$ is equipped with an ideal, which we denote by $H$.
@@
A Hom-coalgebra is a triple $(C,\Delta,\beta)$, where $C$ is a $\mathbb{K}$-vector space, $\Delta:C\rightarrow C\otimes C$ is a linear map with notation $\Delta(c):=\sum_{(c)}c_{(1)}\otimes c_{(2)}$ for all $c\in C$. $\beta:C\rightarrow C$ is a $\mathbb{K}$-linear space homomorphism. And it satisfies
@@
In a similar manner, we will refer to a Hom-coalgebra $(C,\Delta,\beta)$ as $C$. A morphism of Hom-coalgebras is a linear map of the underlying $\mathbb{K}$-linear spaces that commutes with the twisting maps and the comultiplications.
Now, we will give some lemmas and describe their functions. Lemma 2.4 (see [10]) describes the dual structure of the Hom-algebra. Lemma 2.5 (see [11]) introduces the relationship between the morphism of Hom-algebras and its duality. Lemma 2.6 (see [4]) gives the quotient space properties induced by subspaces under linear spaces. Lemma 2.7 (see [11]) explain the fundamental isomorphism theorem of the Hom-algebra.
Lemma 2.4.
Let $(G,\mu,\alpha)$ is a finite-dimensional Hom-algebra, then $(G^{*},\mu^{*},\alpha^{*})$ is a Hom-coalgebra where $\mu^{*}:G^{*}\rightarrow G^{*}\otimes G^{*}$, $\alpha^{*}(a^{*})=a^{*}\alpha$.
Lemma 2.5.
Let $(G,\mu,\alpha)$ and $(G^{{}^{\prime}},\mu^{{}^{\prime}},\alpha^{{}^{\prime}})$ be two finite-dimensional Hom-algebras, then $f:G\rightarrow G^{{}^{\prime}}$ is the morphism of Hom-algebras if and only if $f^{*}:G^{{}^{\prime}*}\rightarrow G^{*}$ is the morphism of Hom-coalgebras.
Lemma 2.6.
Let $A$ and $B$ be two $\mathbb{K}$-linear spaces, $I$ and $J$ be the subspaces of $A$ and $B$ respectively. If there exists a map $\pi:A\otimes B\rightarrow A/I\otimes B/J$, we have that $\operatorname{Ker}(\pi)=A\otimes J+B\otimes I$ and $(A\otimes B)/(A\otimes J+B\otimes I)\cong A/I\otimes B/J$.
Lemma 2.7.
Let $(G,\mu,\alpha)$ and $(G^{{}^{\prime}},\mu^{{}^{\prime}},\alpha^{{}^{\prime}})$ be two Hom-algebras over $\mathbb{K}$, $f:G\rightarrow G^{{}^{\prime}}$ is a morphism of Hom-algebras, $H$ is the ideal of Hom-algebra of $G$, the map $\pi:G\rightarrow G/H$ is a regular full homomorphism. If $\operatorname{Ker}\pi\subseteq\operatorname{Ker}f$, then there exists a unique morphism of Hom-algebras $g:G/H\rightarrow G^{{}^{\prime}}$, so $g\circ\pi=f$.
In order to investigate the structure of Hom-coalgebra induced by the infinite dimensional Hom-algebra and the correspondence of the finite-codimensional ideals in the morphism of Hom-algebras, we will give the definition of the Sweedler duality in Definition 2.8 and the lemma 2.9 below.
@@
The Sweedler duality of $G$ is based on the universal duality $G^{*}$ as follows:
@@
Let $(G,\mu,\alpha)$ and $(G^{{}^{\prime}},\mu^{{}^{\prime}},\alpha^{{}^{\prime}})$ be two Hom-algebras over $\mathbb{K}$, $f:G\rightarrow G^{{}^{\prime}}$ is a morphism of Hom-algebras. If $J$ is a finite-codimensional ideal of $G^{{}^{\prime}}$, then $f^{-1}(J)$ is a finite-codimensional ideal of $G$.
@@
In accordance with the equation (2.3), we obtain the following which holds for any $x\in f^{-1}(J)$ and $a\in G$:
@@
there by establishing that $\alpha(x)\in f^{-1}(J)$. We then proceed with $\alpha(f^{-1}(J))\subseteq f^{-1}(J)$. So $f^{-1}(J)$ is the ideal of Hom-algebra $G$.
@@
Inspired by Liu’ construction [8] of basic algebras, it is sufficient for us to construct the structure of Hom-coalgebras via Sweedler duality now.
@@
$G^{\circ}$ is the subspace of $G^{*}$.
@@
For any $g,h\in G^{\circ}$, let $J$ and $H$ be two finite-codimensional ideals of $G$ such that $\operatorname{Ker}g\supseteq J$ and $\operatorname{Ker}h\supseteq H$. It is clear that $J\cap H$ is an ideal of the Hom-algebra and that $\operatorname{Ker}(g+h)\supseteq J\cap H$.
The following proof explains that $J\cap H$ is finite-codimensional. Since $H/(J\cap H)\cong(J+H)/J\subseteq G/J$, the quotient $H/(J\cap H)$ is a finite-dimensional linear space. Furthermore, since $G/H\cong(G/(J\cap H))/(H/(J\cap H))$, the finite dimension of $G/H$ implies that $dim(G/(J\cap H))=dim(G/H)+dim(H/(J\cap H))$ is also finite. Consequently, $J\cap H$ is a finite-codimensional ideal of $G$. ∎
@@
Let $f\in G^{\circ}$. We show that $\Delta(f)\in G^{\circ}\otimes G^{\circ}$. There exists a finite-codimensional ideal $J$ of $G$ such that $f(J)=0$. Then we do the formal arithmetic
@@
since $\mu(J\otimes G)\subseteq J$ and $\mu(G\otimes J)\subseteq J$. Let a map $\pi\colon G\otimes G\to G\otimes G/(J\otimes G+G\otimes J)$, then we have $\operatorname{Ker}(\pi)=J\otimes G+G\otimes J\subseteq\operatorname{Ker}(f\mu)$. By Lemma 2.7, this induces a unique morphism $\overline{f\mu}\colon G\otimes G/(J\otimes G+G\otimes J)\to\mathbb{K}$ satisfying $\overline{f\mu}\circ\pi=f\circ\mu$. That is to say that $\overline{f\mu}\in(G\otimes G/(J\otimes G+G\otimes J))^{*}$.
@@
Denote $e_{i}^{*}=\overline{e}_{i}^{*}\circ\tilde{\pi}$ Since $\operatorname{Ker}(e_{i}^{*})\supseteq J$, we have $e_{i}^{*}\in G^{\circ}$. Then it is straightforward to see
@@
We therefore have $\operatorname{Ker}(\alpha^{\circ}(g))\supseteq J$, which implies that $\alpha^{\circ}(g)\in G^{\circ}$. So we prove that $\alpha^{\circ}(G^{\circ})\subseteq G^{\circ}$.
Then we will prove that $\Delta\circ\alpha^{\circ}=(\alpha^{\circ}\otimes\alpha^{\circ})\circ\Delta$. Using the equation (2.2), for any $c^{*}\in G^{\circ}$ and $x,y\in G$,
@@
Therefore we have $\Delta\circ\alpha^{\circ}=(\alpha^{\circ}\otimes\alpha^{\circ})\circ\Delta$.
@@
$=\langle c^{*},\alpha(x)yz\rangle$
Similarly, we can get
@@
Due to the Hom-associativity (2.1) of $G$, we have $\alpha(x)yz=xy\alpha(z)$. Thus $(\alpha^{\circ}\otimes\Delta)\circ\Delta=(\Delta\otimes\alpha^{\circ})\circ\Delta$. Therefore we prove this theorem. ∎
In the absence of specific instructions, we will examine the properties of Hom-coalgebra $(G^{\circ},\Delta,\alpha^{\circ})$, the Sweedler duality of $G$ as outlined in Theorem 2. In light of the aforementioned considerations, we proceed to derive the morphism of Hom-coalgebras as follows.
@@
Let $(G,\mu,\alpha)$ and $(G^{{}^{\prime}},\mu^{{}^{\prime}},\alpha^{{}^{\prime}})$ be two Hom-algebras, $(G^{\circ},\Delta,\alpha^{\circ})$ and $(G^{\prime\circ},\Delta^{\prime},\alpha^{\prime\circ})$ two corresponding Sweedler dualities. The map $f:G\rightarrow G^{{}^{\prime}}$ is a morphism of Hom-algebras. Note that the map $f^{\circ}=f^{*}|_{G^{\prime\circ}}:G^{\prime\circ}\rightarrow G^{*}$. Thus, $f^{\circ}$ is a morphism of Hom-coalgebras.
@@
Our primary objective is to demonstrate that $f^{\circ}(G^{\prime\circ})\subseteq G^{\circ}$. For any $b^{*}\in G^{\prime\circ}$, since $J$ is a finite-codimensional ideal of $G^{\prime}$ and is included in $\operatorname{Ker}(b^{*})$, we can utilize Lemma 2.9 to deduce that $f^{-1}(J)$ constitutes a finite-codimensional ideal of the Hom-algebra $G$.
@@
$\langle f^{\circ}(b^{*}),f^{-1}(J)\rangle$ $=\langle b^{*},f\circ f^{-1}(J)\rangle$
$=\langle b^{*},J\rangle$
$=0,$
which shows that $\operatorname{Ker}(f^{\circ}(b^{*}))\supseteq f^{-1}(J)$. In other words, $f^{\circ}(b^{*})\in G^{\circ}$. Consequently, it can be concluded that $f^{\circ}(G^{\prime\circ})\subseteq G^{\circ}$.
@@
Thus, we obtain $\Delta\circ f^{\circ}=(f^{\circ}\otimes f^{\circ})\circ\Delta^{\prime}$.
@@
$\langle\alpha^{\circ}\circ f^{\circ}(b^{*}),a\rangle$ $=\langle b^{*},f\circ\alpha(a)\rangle$
$=\langle b^{*},\alpha^{{}^{\prime}}\circ f(a)\rangle$
$=\langle f^{\circ}\circ\alpha^{\prime\circ}(b^{*}),a\rangle.$

### v2

@@ Abstract.
We establish a dual version of infinite-dimensional Hom-algebras and Hom-modules by using the Sweedler duality construction. Additionally, linear morphisms between infinite-dimensional Hom-algebras (resp. Hom-modules) and Hom-coalgebras (resp. Hom-comodules) are derived under this construction. As an application, we present a Hom-type binary linearly recursive sequences and show that the Sweedler duality construction can be utilized to determine the minimal polynomials of finite-codimensional ideals.
@@
Primary 17A30; Secondary 17A60, 17D30, 05A10
Corresponding author: Chi Zhang
@@ 1. Introduction
Algebraic deformation theory was first studied systematically and in depth by Gerstenhaber [8] and is now one of the most important branches of algebra. In recent years, the introduction of another type of algebraic deformation algebra, namely Hom-algebra, has attracted the attention of many algebraists. In 2006, Hartwig, Larsson, and Silvestrov [11] introduced Hom-Lie algebras to describe $q$-deformations of Witt and Virasoro algebras. The quasi-Lie and quasi-Hom-Lie algebras were introduced by Larsson and Silvestrov [12] in 2005. Later, in 2008, Makhlouf and Silvestrov [17] studied Hom-associative, Hom-Leibniz, Hom-Lie, Lie-admissible, and flexible Hom-algebras. Building on this foundation, the authors [18, 19] generalized the general (co)algebraic structure to Hom-(co)algebras, which essentially extends the concept of associative algebras by deforming the associativity law into a Hom-associativity condition. Later, Yau [26, 27, 28, 30] proposed the concepts of Hom-(co)modules, Hom-Hopf modules and the duality of Hom-algebras, thus enriching the field of Hom-algebra theory. As an emerging field within algebra, (Bi)Hom-algebra is closely related to category theory (see, e.g., [31]), combinatorics [15], braided group representations (see, e.g., [14]), and quantum algebras [2, 16, 20, 21, 22, 23], producing a wealth of research results in recent years.
@@
The exponentials of infinitesimal characters and their convolutions are known to produce characters with elegant combinatorial counterparts [4]. In a Hopf algebra, the set of characters forms a group that acquires a natural structure of an (infinite-dimensional) Lie group under certain conditions, which are often met in practice, such as grading [5]. All these elements relate to the largest dual that can be considered for a Hopf algebra, known as the Sweedler duality. This duality is, in a sense, related to finite-state automaton theory [6] and algebraic combinatorics [24].
@@
In the finite-dimensional case, the structure of the Hom-coalgebra follows directly from the classical duality of the Hom-algebra, and such a conclusion is known. However, this conclusion may not hold in the infinite-dimensional case. This naturally leads to the question of how to get the dual version of such a structure in the infinite-dimensional case. In other words, is it possible to induce the structure of Hom-coalgebras from the general dual of Hom-algebras (resp. from Hom-modules to Hom-comodules) by constructing a correspondence in the infinite-dimensional case? This question was asked by Wang [25] in his conference talk.
@@
In this paper, we will answer his question by investigating the Sweedler duality for Hom-algebras and their induction into Hom-coalgebras in general. We present the new structure of the Hom-coalgebra induced by the Hom-algebra in the context of the Sweedler duality. Furthermore, we extend our results to Hom-modules and establish the connection between Hom-modules and Hom-comodules.
@@
The structure of this paper is as follows. In Section 2, we review some basic concepts and lemmas that discuss the relationship between Hom-algebras and Hom-coalgebras. The structure of Hom-coalgebra under Sweedler duality is given in Theorem 2. A morphism of the Hom-coalgebra is derived using the above condition in Corollary 2.11. In Section 3, we present a right Hom-comodule of Hom-coalgebra, which is induced by the right module of Hom-algebra under Sweedler duality, as shown in Theorem 4. The (right) Hom-comodule morphism is also given in Theorem 5. In the final section 4, we develop a Hom-deformed version of binary linearly recursive sequences and use the Sweedler duality to find the recursive sequences of the minimal polynomial $h(x,y)$.
Notation
Throughout this paper, we assume that the characteristic of the field $\mathbb{K}$ is zero. And we will use the following notation for the mapping role of elements: $f(v):=\langle f,v\rangle$ in which $v$ is any element of some Hom-algebra $G$ and $f$ is an arbitrary element in duality $G^{*}$. Dual maps of Hom-algebra $(G,\mu,\alpha)$ are $\mu^{*}\!:G^{*}\!\to G^{*}\!\otimes\!G^{*}$, $\langle\mu^{*}(f),x\!\otimes\!y\rangle=\langle f,\mu(x\!\otimes\!y)\rangle$, $\alpha^{*}(f)=f\!\circ\!\alpha$. And for a right Hom-module $(M,\psi,\gamma)$ over $G$ we set $\psi^{*}\!:M^{*}\!\to M^{*}\!\otimes\!G^{*}$, $\langle\psi^{*}(\xi),m\!\otimes\!g\rangle=\langle\xi,\psi(m\!\otimes\!g)\rangle$, $\gamma^{*}(\xi)=\xi\!\circ\!\gamma$. For more details on notation, this paper follows the setting of [8].
2. Sweedler duality for Hom-algebras
@@
In this section, we will first give an overview of the basic concepts related to the structures of Hom-(co)algebras and the interrelationship between different Hom-(co)algebras (see Definition 2.3 in [29]). We then move on to introduce the definition of the Sweedler duality, which differs from the conventional duality (see Chapter 6 in [8]).
@@
A Hom-associative algebra is a triple $(G,\mu,\alpha)$, where $G$ is a linear space, $\mu:G\otimes G\rightarrow G$ is a bilinear multiplication with $\mu(g\otimes g^{\prime})=gg^{\prime}$. For simplicity, when there is no confusion, we will omit the notation $\mu$. $\alpha:G\rightarrow G$ is a homomorphism. Otherwise, it satisfies the following conditions
@@
In the following of this paper, we will abbreviate Hom-associative algebra to Hom-algebra, and sometimes we refer to a Hom-algebra $(G,\mu,\alpha)$ as $G$. Now, for two Hom-algebras $(G,\mu,\alpha)$ and $(G^{\prime},\mu^{\prime},\alpha^{\prime})$, let $f:G\to G^{\prime}$ be a linear map, then $f$ is said to be a morphism of Hom-algebras if it satisfies the following conditions
@@
Let $(G,\mu,\alpha)$ be a Hom-algebra, then a subspace $H$ of $G$ is a Hom-associative subalgebra of $G$ if $bc\in H$ and $\alpha(b)\in H$ for any $b,c\in H$. In particular, if both $ab\in H$ and $ba\in H$ hold for any $a\in G$ and $b\in H$, we say that $H$ is an ideal of $G$.
@@
A Hom-coalgebra is a triple $(C,\Delta,\beta)$, where $C$ is a linear space, $\Delta:C\rightarrow C\otimes C$ is a linear map with notation $\Delta(c):=\sum_{(c)}c_{(1)}\otimes c_{(2)}$ for all $c\in C$. And $\beta:C\rightarrow C$ is a homomorphism which satisfies
@@
Similarly, we will denote a Hom-coalgebra $(C,\Delta,\beta)$ by $C$. A morphism of Hom-coalgebras is a linear map of the underlying $\mathbb{K}$-linear spaces that commutes with the twisting maps and the comultiplications.
Now we give some lemmas that describe some fundamental properties of Hom-(co)algebras.
Lemma 2.4 (see [18]).
Let $(G,\mu,\alpha)$ be a finite-dimensional Hom-algebra, then $(G^{*},\mu^{*},\alpha^{*})$ is a Hom-coalgebra where $\mu^{*}:G^{*}\rightarrow G^{*}\otimes G^{*}$, $\alpha^{*}(a^{*})=a^{*}\alpha$.
Lemma 2.5 (see [18]).
Let $(G,\mu,\alpha)$ and $(G^{\prime},\mu^{\prime},\alpha^{\prime})$ be two finite-dimensional Hom-algebras, then $f:G\rightarrow G^{\prime}$ is a morphism of Hom-algebras if and only if $f^{*}:G^{\prime*}\rightarrow G^{*}$ is a morphism of Hom-coalgebras.
Lemma 2.6 (see [7]).
Let $A$ and $B$ be two linear spaces, $I$ and $J$ be two subspaces of $A$ and $B$ respectively. If there exists a map $\pi:A\otimes B\rightarrow A/I\otimes B/J$, we have $\operatorname{Ker}(\pi)=A\otimes J+B\otimes I$ and $(A\otimes B)/(A\otimes J+B\otimes I)\cong A/I\otimes B/J$.
Lemma 2.7 (see [18]).
Let $(G,\mu,\alpha)$ and $(G^{\prime},\mu^{\prime},\alpha^{\prime})$ be two Hom-algebras, the map $f:G\rightarrow G^{\prime}$ be a morphism of Hom-algebras, $H$ be the ideal of Hom-algebra of $G$, the map $\pi:G\rightarrow G/H$ be a regular full homomorphism. If $\operatorname{Ker}\pi\subseteq\operatorname{Ker}f$, there exists a unique Hom-algebra morphism $g:G/H\rightarrow G^{{}^{\prime}}$, which satisfies $g\circ\pi=f$.
Remark 2.8.
A regular full homomorphism means the canonical quotient map $\pi:G\rightarrow G/I$ associated with an ideal $I\subseteq G$; that is, a homomorphism of Hom-associative algebras which is surjective (hence a regular epimorphism) and whose kernel equals $I$.
In order to study the structure of the Hom-coalgebra induced by the infinite-dimensional Hom-algebra and the correspondence of the finite-codimensional ideals in the morphism of Hom-algebras, we will give the definition of Sweedler duality in Definition 2.9 and the lemma 2.10 below.
@@
Let $G$ be a Hom-algebra over $\mathbb{K}$. The Sweedler duality of $G$ is based on the classical algebraic duality $G^{*}$ as follows:
@@
Let $(G,\mu,\alpha)$ and $(G^{\prime},\mu^{\prime},\alpha^{\prime})$ be two Hom-algebras, $f:G\rightarrow G^{\prime}$ be a morphism of Hom-algebras. If $J$ is a finite-codimensional ideal of $G^{\prime}$, then its complete inversion $f^{-1}(J)$ is also a finite-codimensional ideal of $G$.
@@
According to equation (2.3), we obtain the following, which holds for any $x\in f^{-1}(J)$ and $a\in G$:
@@
there by establishing $\alpha(x)\in f^{-1}(J)$. We then proceed with $\alpha(f^{-1}(J))\subseteq f^{-1}(J)$. So $f^{-1}(J)$ is the ideal of Hom-algebra $G$.
@@
Inspired by Liu’s construction of classical algebras [13], it is sufficient for us to construct the structure of Hom-coalgebras via the Sweedler duality now.
@@
Let $G$ be a Hom-algebra. Then $G^{\circ}$ is the subspace of $G^{*}$.
@@
For any $g,h\in G^{\circ}$, let $J$ and $H$ be two finite-codimensional ideals of $G$ such that $\operatorname{Ker}g\supseteq J$ and $\operatorname{Ker}h\supseteq H$. It is clear that $J\cap H$ is an ideal of the Hom-algebra and $\operatorname{Ker}(g+h)\supseteq J\cap H$.
Our next step is to explain $J\cap H$ is finite-codimensional. Since $H/(J\cap H)\cong(J+H)/J\subseteq G/J$, the quotient $H/(J\cap H)$ is a finite-dimensional linear space. Furthermore, since $G/H\cong(G/(J\cap H))/(H/(J\cap H))$, the finite dimension of $G/H$ implies that $dim(G/(J\cap H))=dim(G/H)+dim(H/(J\cap H))$ is also finite. Consequently, $J\cap H$ is a finite-codimensional ideal of $G$. ∎
@@
Let $f\in G^{\circ}$, we want to show that $\Delta(f)\in G^{\circ}\otimes G^{\circ}$. There exists a finite-codimensional ideal $J$ of $G$ such that $f(J)=0$. Then we do the formal arithmetic
@@
since $\mu(J\otimes G)\subseteq J$ and $\mu(G\otimes J)\subseteq J$. Let a map $\pi\colon G\otimes G\to G\otimes G/(J\otimes G+G\otimes J)$, then we have $\operatorname{Ker}(\pi)=J\otimes G+G\otimes J\subseteq\operatorname{Ker}(f\mu)$. By Lemma 2.7, this induces a unique morphism $\overline{f\mu}\colon G\otimes G/(J\otimes G+G\otimes J)\to\mathbb{K}$ satisfying $\overline{f\mu}\circ\pi=f\circ\mu$. That is to say $\overline{f\mu}\in(G\otimes G/(J\otimes G+G\otimes J))^{*}$.
@@
Denote $e_{i}^{*}=\overline{e}_{i}^{*}\circ\tilde{\pi}$. Because of $\operatorname{Ker}(e_{i}^{*})\supseteq J$, we have $e_{i}^{*}\in G^{\circ}$. It is straightforward that
@@
Hence $\operatorname{Ker}(\alpha^{\circ}(g))\supseteq J$ follows, which implies that $\alpha^{\circ}(g)\in G^{\circ}$.
Using the equation (2.2), for any $c^{*}\in G^{\circ}$ and $x,y\in G$, we have
@@
Equivalently speaking, we have $\Delta\circ\alpha^{\circ}=(\alpha^{\circ}\otimes\alpha^{\circ})\circ\Delta$.
@@
$=\langle c^{*},\alpha(x)yz\rangle.$
Similarly,
@@
Due to the Hom-associativity (2.1) of $G$, we have $\alpha(x)yz=xy\alpha(z)$. Thus $(\alpha^{\circ}\otimes\Delta)\circ\Delta=(\Delta\otimes\alpha^{\circ})\circ\Delta$. ∎
Without specific instructions, we will examine the properties of the Hom-coalgebra $(G^{\circ},\Delta,\alpha^{\circ})$, Sweedler duality of $G$ as described in Theorem 2. Based on these considerations, we then derive the morphism of Hom-coalgebras accordingly.
@@
Let $(G,\mu,\alpha)$ and $(G^{\prime},\mu^{\prime},\alpha^{\prime})$ be two Hom-algebras, $(G^{\circ},\Delta,\alpha^{\circ})$ and $(G^{\prime\circ},\Delta^{\prime},\alpha^{\prime\circ})$ two corresponding objects under the Sweedler duality. The map $f:G\rightarrow G^{\prime}$ is a morphism of Hom-algebras. Note that the map $f^{\circ}=f^{*}|_{G^{\prime\circ}}:G^{\prime\circ}\rightarrow G^{*}$. Thus, $f^{\circ}$ is a morphism of Hom-coalgebras.
@@
Our primary goal is to show that $f^{\circ}(G^{\prime\circ})\subseteq G^{\circ}$. For any $b^{*}\in G^{\prime\circ}$, since $J$ is a finite-codimensional ideal of $G^{\prime}$ and lies in $\operatorname{Ker}(b^{*})$, by using Lemma 2.10, we can deduce that $f^{-1}(J)$ constitutes a finite-codimensional ideal of the Hom-algebra $G$.
@@
$\langle f^{\circ}(b^{*}),f^{-1}(J)\rangle=\langle b^{*},f\circ f^{-1}(J)\rangle=\langle b^{*},J\rangle=0,$
which shows that $\operatorname{Ker}(f^{\circ}(b^{*}))\supseteq f^{-1}(J)$. In other words, $f^{\circ}(b^{*})\in G^{\circ}$. Consequently, $f^{\circ}(G^{\prime\circ})\subseteq G^{\circ}$.
@@
It means we obtain that $\Delta\circ f^{\circ}=(f^{\circ}\otimes f^{\circ})\circ\Delta^{\prime}$.
@@
$\langle\alpha^{\circ}\circ f^{\circ}(b^{*}),a\rangle=\langle b^{*},f\circ\alpha(a)\rangle=\langle b^{*},\alpha^{\prime}\circ f(a)\rangle=\langle f^{\circ}\circ\alpha^{\prime\circ}(b^{*}),a\rangle.$